# Beyond Markovian: Reflective Exploration via Bayes-Adaptive RL for LLM Reasoning

### v1

@@ Abstract
Large Language Models (LLMs) trained via Reinforcement Learning (RL) have exhibited strong reasoning capabilities and emergent reflective behaviors, such as backtracking and error correction. However, conventional Markovian RL confines exploration to the training phase to learn an optimal deterministic policy and depends on the history contexts only through the current state. Therefore, it remains unclear whether reflective reasoning will emerge during Markovian RL training, or why they are beneficial at test time. To remedy this, we recast reflective exploration within the Bayes-Adaptive RL framework, which explicitly optimizes the expected return under a posterior distribution over Markov decision processes. This Bayesian formulation inherently incentivizes both reward-maximizing exploitation and information-gathering exploration via belief updates. Our resulting algorithm, BARL, instructs the LLM to stitch and switch strategies based on the observed outcomes, offering principled guidance on when and how the model should reflectively explore. Empirical results on both synthetic and mathematical reasoning tasks demonstrate that BARL outperforms standard Markovian RL approaches at test time, achieving superior token efficiency with improved exploration effectiveness. Our code is available at https://github.com/shenao-zhang/BARL.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive reasoning abilities, such as in solving complex math problems. A key factor driving this progress is the use of Chain-of-Thought (CoT) reasoning (wei2022chain, ), where the model engages in intermediate deliberation before producing an answer. Building on this, recent advances have employed Reinforcement Learning (RL) to further enhance LLM reasoning by optimizing for verifiable outcome rewards (jaech2024openai, ; guo2025deepseek, ; yu2025dapo, ; wei2025swe, ). Notably, RL-trained models have exhibited emergent behaviors such as generating long CoTs and engaging in self-reflection, a process of backtracking to previous states to correct earlier mistakes, also known as the “Aha moment” (guo2025deepseek, ; zeng2025simplerl, ). However, despite these compelling phenomena, it remains unclear why and under what conditions reflective reasoning is beneficial at test time, or whether such behaviors will emerge through conventional RL training.
Prevalent views attempt to explain the usefulness of test-time reflections as exploratory steps that provide additional contexts for more optimal decision-making. Yet in Markovian RL, the exploration–exploitation trade-off is resolved entirely during training: the agent interleaves exploration and exploitation to learn a training-time optimal deterministic policy, but switches to pure exploitation at test time. As a result, conventional RL allows a Markovian policy to be optimal by simply memorizing, i.e., deterministically outputting, training solutions once they are encountered by stochastic exploratory policies through repeated trial-and-error. Moreover, the Markov assumption restricts the policy to condition decisions solely on the current state rather than on contextual information gathered through exploration, offering no incentives to adaptively explore with reflections. In summary, under conventional Markovian RL, there is no guarantee that reflective explorations will emerge during training, nor does it explain why such explorations might be advantageous at test time.
@@
To address this gap, we propose grounding reflective reasoning with Bayes-Adaptive RL, which explicitly optimizes for test-time generalization by maximizing the expected return under a posterior distribution over Markov Decision Processes (MDPs). The objective incentivizes both reward-seeking actions and epistemic explorations that gather information to reduce the MDP’s uncertainty, such as the uncertainty regarding the progress made by different actions. This enables the model to adapt on-the-fly at test time by updating its beliefs and switching strategies based on observed outcomes, naturally giving rise to reflective exploration behaviors. We prove that the expected return of an adaptive policy can be exponentially higher than the optimal Markovian policy at test time.
@@
Building upon this formulation, we introduce a novel algorithm, Bayes-Adaptive RL for LLM Reasoning (BARL). For each prompt, BARL performs online rollouts to generate a set of candidate answers, each associated with an MDP hypothesis. The state-action value is then computed by weighting each hypothesis according to the model’s current belief, with penalties applied for mismatches between predicted and observed rewards, thereby signaling when to switch strategies. BARL provides a principled mechanism for integrating and revising plausible strategies, analogous to linearizing best-of-N reasoning, but with explicit step-level guidance on when and how the model should reflectively explore.
@@
To illustrate the benefits of BARL, we begin with a synthetic task designed to mirror test-time generalization in LLM reasoning. The agent receives a reward only when it repeats a prompt token exactly three times, but the training and testing prompt tokens differ. Standard Markovian RL memorizes the training solutions and fails to generalize. In contrast, BARL learns to switch strategies by eliminating hypotheses, ultimately discovering the ground-truth MDP for optimal behavior.
@@
We further evaluate BARL on math reasoning tasks using various LLMs, including Qwen2.5-Math-1.5B, Qwen2.5-Math-7B, and R1-Distill-Llama-8B. Across these models, BARL consistently outperforms Markovian RL algorithms, such as GRPO and a strong progress-reward baseline, on multiple benchmarks. BARL achieves significantly greater token efficiency, requiring up to 39% fewer average tokens than the progress baseline, 50% fewer than GRPO, and over 90% fewer than the Qwen2.5-Math-1.5B base model. Moreover, we observe no strong correlation between overall model performance and the frequency of reflections. Instead, BARL’s advantage stems from more efficient exploration and more effective thinking tokens. We summarize the key takeaways of this paper as follows:
@@
Key Takeaways: Why, How, and When Should LLMs Reason Reflectively Why: Markovian RL neither ensures the emergence of reflective exploration nor explains its benefits at test time since (1) exploration is confined only to the training phase to learn, and purely exploit at test time, an optimal deterministic policy that can merely memorize training solutions, and (2) the state-conditional policy lacks incentives to collect additional contexts and backtrack. In contrast, Bayes-Adaptive RL, by optimizing test-time generalizability, encourages explorations to gather contextual information that reduces the MDP uncertainty. How: BARL provides a principled way to stitch plausible strategies by maintaining a posterior over MDP hypotheses, each associated with a sampled candidate answer. Reflective exploration emerges naturally through hypothesis elimination, enabling on-the-fly adaptation. When: LLMs should self-reflect when discrepancies arise between their internal beliefs and cumulative reward feedback—signaling strategy switching by downweighting hypotheses that have high belief probabilities but are unlikely to be optimal given previous observations.
@@ 2 Related Work
As an emerging capability of model scale, LLMs can generate intermediate CoTs to solve complex reasoning tasks (wei2022chain, ; kojima2022large, ) and scale test-time performance by allocating more thinking tokens (snell2024scaling, ; brown2024large, ). Early efforts enhanced LLM reasoning via supervised fine-tuning on human-annotated data (cobbe2021training, ; yue2023mammoth, ; yu2023metamath, ) or linearized search traces (lehnert2024beyond, ; gandhi2024stream, ). However, due to the distribution shift between LLM responses and curated data, LLM-generated data has proven effective through rejection sampling (dong2023raft, ; yuan2023scaling, ) by filtering out low-quality rationales (zelikman2022star, ; zelikman2024quiet, ) or with EM iterations (singh2023beyond, ; zhong2025brite, ). Recently, RL has gained increasing interest for improving reasoning (aksitov2023rest, ; havrilla2024teaching, ; wang2024offline, ; shao2024deepseekmath, ). Process rewards (uesato2022solving, ; lightman2023let, ) with Monte Carlo unrolls (kazemnejad2024vineppo, ; wang2023math, ; wang2024multi, ; luo2024improve, ) offer finer-grained feedback but are computationally expensive. Outcome-reward RL (guo2025deepseek, ) demonstrates emergent deliberative reasoning abilities such as self-reflection. Yet, limited work has investigated the underlying mechanisms of such behaviors. In fact, recent findings suggest that reflections do not consistently emerge from RL training and exhibit weak correlation with performance (liu2025understanding, ). Similar to our work, xiang2025towards ; qu2025optimizing also study the generalization of LLMs, from a meta-RL (duan2016rl, ) perspective: xiang2025towards justify deliberative reasoning as providing extra contexts, and qu2025optimizing use progress reward (setlur2024rewarding, ) to reduce regret in outcome-reward RL. Our method differs from xiang2025towards in that we ground reflective reasoning in environment rewards, rather than relying solely on the internal CoT states generated by the model itself. Compared to qu2025optimizing , which rewards golden strategies that make progress towards the correct answer, BARL additionally encourages exploring plausible strategies under the Bayesian framework, allowing it to account for uncertainty during both training and testing. We experimentally compare against a variant of qu2025optimizing that estimates progress using answer probability differences instead of more expensive Monte Carlo unrolls. Besides, unlike wang2023hypothesis ; qiu2023phenomenal that manually design hypothesis proposal–selection pipelines, our method achieves this through a more principled RL optimization procedure.
@@
Conventional RL explores only during training, e.g. via $\epsilon$-greedy noise, and exploits the optimal deterministic policy when deployed. Exceptions include works that explicitly optimize maximum entropy objectives (haarnoja2018soft, ) to learn stochastic policies, primarily to accelerate training convergence in settings where evaluation remains in-distribution, such as robotic control. Bayes-Adaptive RL (bellman1959adaptive, ; duff2002optimal, ; guez2012efficient, ; ghavamzadeh2015bayesian, ; lidayan2024bamdp, ; zhang2024self, ; liu2023reason, ) has been studied to pursue the optimal exploration-exploitation trade-off in uncertain environments to improve generalizability. When the true MDP identity is latent and must be inferred from interaction (states, actions, rewards), Bayesian RL connects naturally to Partially Observable MDPs (duff2001monte, ; ghosh2021generalization, ). Exact solutions of the Bayesian RL objective are often intractable, prompting the development of approximate methods (guez2012efficient, ; arumugamreducing, ; chen2024bayes, ). In our work, we adopt policy gradient that operates over candidate answers, which differs from ghosh2022offline that leverage value ensembles in offline RL and qiu2025bayesian that applies SFT on an oracle Bayesian model’s outputs.
@@ 3 Problem Formulation
LLM Reasoning via RL.
To enable the LLM policy $\pi_{\theta}$ to reason, we first consider the finite-horizon MDP defined by the state space $\mathcal{S}$, action space $\mathcal{A}$, horizon $T$, and reward function $r(s,a)$, where $s\in\mathcal{S}$ and $a\in\mathcal{A}$. Here, the initial state $s_{0}$ is the prompt, and the action $a_{t}$ is the $t$-th step of the CoT, which can be either separated by special tokens (wang2023math, ) or defined as a fixed length of reasoning tokens (luo2024improve, ). We adopt the latter definition due to its simplicity. The state transition is deterministic by appending the new reasoning step, i.e., $s_{t+1}=s_{t}+a_{t}$. Prior work (uesato2022solving, ; guo2025deepseek, ) employs an outcome-level reward $\text{verifier}(s_{T},y_{s_{0}}^{*})$, which uses a verifier to perform a regular expression match (either 00 or $1$) between $s_{T}$ and the ground-truth answer $y_{s_{0}}^{*}$ corresponding to the prompt $s_{0}$. We extend this sparse-reward setting by incorporating a progress reward (setlur2024rewarding, ; qu2025optimizing, ), which quantifies the increase of the model’s probability of outputting $y_{s_{0}}^{*}$ after appending $a$ at the CoT $s$, i.e., for $10\leq t\leq T-$ 1:
@@
[…]
For the Markovian RL objective $\mathcal{J}_{\text{Markov}}(\pi_{\theta}):=\mathbb{E}_{s_{0},\pi_{\theta}}[\sum_{t=0}^{T-1}r(s_{t},a_{t})+\text{verifier}(s_{T},y_{s_{0}}^{*})]$, this reward definition allows us to use telescoping in a way similar to reward shaping (ng1999policy, ) to obtain
$\mathop{\mathrm{argmax}}_{\pi_{\theta}}\mathcal{J}_{\text{Markov}}(\pi_{\theta})=\mathop{\mathrm{argmax}}_{\pi_{\theta}}\mathbb{E}_{s_{0},\pi_{\theta}}\bigl{[}\pi_{\theta}(y_{s_{0}}^{*}\mid s_{0}+a_{0:T-1}+\text{</think>})+\text{verifier}(s_{T+1},y_{s_{0}}^{*})\bigr{]},$
i.e., the optimal Markovian policy generates $a_{0:T-1}$ to maximize the likelihood of the ground-truth $y_{s_{0}}^{*}$ and its verifier-evaluated correctness. The gradients for Markovian policies are
$\nabla_{\theta}\mathcal{J}_{\text{Markov}}(\pi_{\theta})=\mathbb{E}_{s_{0},\pi_{\theta}}\biggl{[}\sum_{t=0}^{T-1}\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})\cdot Q^{\pi_{\theta}}(s_{t},a_{t})\biggr{]},$ (3.2)
where $Q^{\pi_{\theta}}$ is the state-action value or advantage function (schulman2015high, ). The above setups consider the case when the environment is predefined with certainty. The definitions naturally extend to any MDP $\mathcal{M}:=(\mathcal{S},\mathcal{A},r_{\mathcal{M}},T)$ where $r_{\mathcal{M}}$ is defined w.r.t. the answer $y_{s_{0}}^{\mathcal{M}}$. The Q-value is then
$Q^{\pi_{\theta}}_{\mathcal{M}}(s_{t},a_{t})$ $=\mathbb{E}_{\pi_{\theta}}\biggl{[}\sum_{t^{\prime}=t}^{T-1}r_{\mathcal{M}}(s_{t^{\prime}},a_{t^{\prime}})+\text{verifier}(s_{T},y_{s_{0}}^{*})\biggr{]}$ (3.3)
$=\mathbb{E}_{\pi_{\theta}}\bigl{[}\pi_{\theta}(y_{s_{0}}^{\mathcal{M}}\mid s_{t}+a_{t:T-1}+\text{</think>})-\pi_{\theta}(y_{s_{0}}^{\mathcal{M}}\mid s_{t}+\text{</think>})+\text{verifier}(s_{T},y_{s_{0}}^{*})\bigr{]}.$
Figure 1: An example of reflective reasoning.
@@
We define reflective exploration as the pattern in which the LLM backtracks to a prior state after an exploratory step to take different actions at that state. Specifically, a natural language reflective reasoning step such as “Let’s reconsider the geometric relationship” corresponds to a backtracking action that semantically disregards the previous one or more steps. We illustrate this using a binary search tree as in the right figure: for the trajectory $s_{0}\shortrightarrow s_{1}\shortrightarrow s_{2}\shortrightarrow s_{1}\shortrightarrow s_{3}$, $s_{1}\shortrightarrow s_{2}$ is an exploration step, $s_{2}\shortrightarrow s_{1}$ is a reflective step that signals the strategy switch from $s_{2}$ to $s_{3}$, and the “geometric relationship” in the above example originates from $s_{1}$.
4 The Necessity of Bayes-Adaptive RL for Reflective Reasoning
Markovian RL.
When the underlying MDP is known with certainty, the Markov property ensures that the policy and value depend on the history $h_{t}=(s_{0},a_{0},r_{0},\ldots,s_{t})$ only through the state $s_{t}$, i.e., $Q^{\pi}(h_{t},a_{t})=Q^{\pi}(s_{t},a_{t})$. In this setting, exploratory actions that aim to enrich the history $h_{t}$ with additional contexts, such as incorrect attempts followed by backtracking, are unnecessary, as the current state $s_{t}$ already encodes all relevant information for optimal decision-making [^1] Even if we view the state as the CoT history, i.e., explorations such as $s_{1}\shortrightarrow s_{2}$ are not semantically disregarded after backtracking to $s_{1}$, it still differs from $h_{t}$ where reward histories are additionally contained..
Moreover, the optimal Q-function is $Q^{*}=\max_{\pi}Q^{\pi}$ and the optimal policy $\pi^{*}$ is greedy w.r.t. $Q^{*}$. That is, $\pi^{*}$ is a deterministic policy, where $\pi^{*}(a^{\prime}|s)=1$ for $a^{\prime}=\mathop{\mathrm{argmax}}_{a}Q^{*}(s,a)$.
Theorem 4.1.
Optimality of Markovian RL is attained by deterministic non-reflective policies.
Reflective policies are more suboptimal than non-reflective policies in both discounted infinite-horizon and finite-horizon MDPs, as the $Q$-value of the wrong action is no larger than that of the correct action since more tokens are needed to correct the error. Explorations should only occur during training, in a trial-and-error manner with repeated episodes, to discover the golden answers. The Markovian RL objective allows the optimal policy that memorizes these training answers to be fully exploited, with no incentive to adaptively explore with reflections.
For the non-standard undiscounted infinite-horizon MDPs, where LLMs are encouraged to generate infinite tokens without concerning token efficiency, reflective policies may be as optimal as non-reflective ones. This is because the $Q$-value of the wrong action can match that of the correct one if the error is eventually corrected through reflection. This observation provides a partial explanation for the emergence of “Aha moment” with long CoT. However, even in such settings, reflective reasoning may still fail to emerge under Markovian RL, particularly if golden answers are discovered either directly or by pruning incorrect exploratory steps. In other words, this only explains why reflective explorations can appear during Markovian RL, instead of why these behaviors are preferable to simply memorizing training solutions, nor whether they will emerge during training.
Next, we present Bayes-Adaptive RL, which explicitly optimizes for test-time generalization and naturally induces reflective explorations.
Bayes-Adaptive RL.

### v2

@@ Abstract
Large Language Models (LLMs) trained via Reinforcement Learning (RL) have exhibited strong reasoning capabilities and emergent reflective behaviors, such as rethinking and error correction, as a form of in-context exploration. However, the Markovian policy obtained from conventional RL training does not give rise to reflective exploration behaviors since the policy depends on the history only through the state and therefore has no incentive to enrich identical states with additional context. Instead, RL exploration is only useful during training to learn the optimal policy in a trial-and-error manner. Therefore, it remains unclear whether reflective reasoning will emerge during RL, or why it is beneficial. To remedy this, we recast reflective exploration within a Bayesian RL framework, which optimizes the expected return under a posterior distribution over Markov decision processes induced by the training data. This Bayesian formulation admits uncertainty-adaptive policies that, through belief updates, naturally incentivize information-gathering actions and induce self-reflection behaviors. Our resulting algorithm, BARL, instructs the LLM to stitch and switch strategies based on the observed outcomes, offering principled guidance on when and how the model should reflectively explore. Empirical results on both synthetic and mathematical reasoning tasks demonstrate that BARL outperforms conventional RL approaches, achieving superior test-time performance and token efficiency. Our code is available at https://github.com/shenao-zhang/BARL.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive reasoning abilities, such as in solving complex math problems. A key factor driving this progress is the use of Chain-of-Thought (CoT) reasoning (wei2022chain), where the model engages in intermediate deliberation before producing an answer. Building on this, recent advances have employed Reinforcement Learning (RL) to further enhance LLM reasoning by optimizing for verifiable outcome rewards (jaech2024openai; guo2025deepseek; yu2025dapo; wei2025swe). Notably, RL-trained models have exhibited emergent behaviors such as engaging in self-reflection, a process of revisiting previous states to correct earlier mistakes (guo2025deepseek; zeng2025simplerl), which offers a potential way for test-time scaling (snell2024scaling; brown2024large) by generating longer CoTs. However, despite these compelling phenomena, it remains unclear why and under what conditions self-reflections are beneficial, or whether such behaviors will emerge through conventional RL training.
In this work, we use reflective exploration to refer to behavior where the model revisits a previously visited state and, informed by the additional context it has acquired, alters its subsequent actions. This definition links self-reflective directly to exploration under uncertainty, providing a principled basis for when and why strategy switches should occur rather than treating them as ad hoc quirks of the model’s CoTs. We formalize this via strictly uncertainty-adaptive policies whose actions depend nontrivially on the belief. Since conventional RL typically admits a Markovian optimal policy, reflective exploration is not induced since the policy depends on the history only through the state and is thus not incentivized to enrich identical states with additional context. Instead, RL explores during training through repeated trial and error, with no incentives for further exploration when parameters are frozen after deployment. Therefore, under conventional RL, there is no guarantee that self-reflections will emerge during training, nor does it explain why they might be advantageous when the trained policy is deployed at test time.
@@
To address this gap, we ground reflective exploration with Bayesian RL, which maximizes the expected return in the deployment stage under a posterior distribution over Markov Decision Processes (MDPs) induced by the training data. The objective incentivizes both reward-seeking actions and exploration actions that gather information to reduce the MDP’s uncertainty, such as the reward uncertainty regarding the progress made by different actions. We show that optimizing this Bayesian RL objective yields policies that adapt on-the-fly by updating their beliefs and switching strategies based on observed outcomes, naturally leading to reflective exploration behaviors. Moreover, we prove that the optimal adaptive policy can achieve arbitrarily higher Bayes-expected return than any optimal Markovian policy.
@@
Building upon this formulation, we introduce a novel algorithm, Bayes-Adaptive RL for LLM Reasoning (BARL). For each prompt, BARL performs online rollouts to generate a set of candidate answers, each associated with an MDP hypothesis. The state-action value is then computed by weighting each hypothesis according to the model’s current belief, with penalties applied for mismatches between predicted and observed rewards, thereby signaling when to switch strategies. BARL provides a principled mechanism for integrating and revising plausible strategies, analogous to linearizing best-of-N, but with explicit guidance on when and how the model should reflectively explore.
@@
To illustrate the benefits of BARL, we begin with a synthetic task designed to mirror the training-deployment procedure in LLM reasoning. The agent receives a reward only when it repeats a prompt token exactly three times, but the training and evaluation prompt tokens differ. Conventional RL fails to generalize beyond the training prompt tokens. In contrast, BARL learns to switch strategies by eliminating hypotheses, ultimately discovering the ground-truth MDP for optimal behavior.
@@
We further evaluate BARL on math reasoning tasks using various LLMs, including Qwen2.5-Math-1.5B, Qwen2.5-Math-7B, and R1-Distill-Llama-8B. Across these models, BARL consistently outperforms conventional RL algorithms, such as GRPO and a strong progress-reward baseline, on multiple benchmarks. BARL achieves significantly greater token efficiency, requiring up to 1.63x fewer average tokens than the progress baseline, 2x fewer than GRPO, and over 10x fewer than the Qwen2.5-Math-1.5B base model. Moreover, we observe no strong correlation between overall performance and the frequency of reflections. Instead, BARL’s advantage stems from more efficient exploration and more effective thinking tokens.
@@
Key Takeaways: Why, How, and When Should LLMs Reflectively Explore Why: RL neither ensures the emergence of reflective exploration nor explains its benefits since (1) RL attains Markovian optimal policies, which cannot exhibit reflective exploration that collects additional context and act differently at the same state, and (2) RL exploration is confined to trial-and-error during training, leaving no mechanism that encourages adaptive exploration at test time. In contrast, Bayesian RL, by optimizing Bayes-expected return during deployment, encourages exploration to gather contextual information that reduces the MDP uncertainty. How: BARL provides a principled way to stitch plausible strategies by maintaining a posterior over MDP hypotheses, each associated with a sampled candidate answer. Reflective exploration emerges through hypothesis elimination, enabling on-the-fly adaptation. When: LLMs should self-reflect when discrepancies arise between their internal beliefs and cumulative reward feedback—signaling strategy switching by downweighting hypotheses that have high belief probabilities but are unlikely to be optimal given previous observations.
@@ 2 Related Work
As an emerging capability of model scale, LLMs can generate intermediate CoTs to solve complex reasoning tasks (wei2022chain; kojima2022large) and scale test-time performance by allocating more thinking tokens (snell2024scaling; brown2024large). Early efforts enhanced LLM reasoning via supervised fine-tuning on human-annotated data (cobbe2021training; yue2023mammoth; yu2023metamath) or linearized search traces (lehnert2024beyond; gandhi2024stream). However, due to the distribution shift between LLM responses and curated data, LLM-generated data has proven effective through rejection sampling (dong2023raft; yuan2023scaling) by filtering out low-quality rationales (zelikman2022star; zelikman2024quiet) or with EM iterations (singh2023beyond; zhong2025brite). Recently, RL has gained increasing interest for improving reasoning (aksitov2023rest; havrilla2024teaching; wang2024offline; shao2024deepseekmath). Process rewards (uesato2022solving; lightman2023let) with Monte Carlo unrolls (kazemnejad2024vineppo; wang2023math; wang2024multi; luo2024improve) offer finer-grained feedback but are computationally expensive. Outcome-reward RL (guo2025deepseek) demonstrates emergent deliberative reasoning abilities such as self-reflection. Yet, limited work has investigated the underlying mechanisms of such behaviors. In fact, recent findings suggest that reflections do not consistently emerge from RL training and exhibit weak correlation with performance (liu2025understanding). Similar to our work, xiang2025towards; qu2025optimizing also study reflective exploration of LLMs, from a meta-RL (duan2016rl) perspective: xiang2025towards justifies deliberative reasoning as providing extra CoT state contexts, and qu2025optimizing uses progress reward (setlur2024rewarding) to reduce regret in outcome-reward RL. Our method differs from xiang2025towards in that we ground reflective reasoning in environment rewards, rather than relying solely on the internal CoT states generated by the model itself. Compared to qu2025optimizing, which rewards strategies that make progress towards the correct answer, BARL additionally encourages exploring plausible strategies under the Bayesian framework. We defer a more detailed discussion to the end of Section 4. We experimentally compare with a variant of qu2025optimizing that estimates progress using answer probability differences. Prior work has also brought Bayesian exploration principles to LLMs: arumugam2025toward implements Posterior Sampling for Reinforcement Learning (PSRL), a statistically-efficient algorithm, using LLMs as modular components, while dwaracherla2024efficient shows that active query selection with epistemic neural networks and double Thompson sampling reduces the amount of human feedback required to train reward models for RLHF. In contrast, we optimize a Bayesian objective directly during RL fine-tuning of the LLM policy, rather than at the level of external algorithmic scaffolding or data collection, to provide a step-level guidance on when and how LLMs should self-reflect. Besides, our method differs from wang2023hypothesis; qiu2023phenomenal that manually designs hypothesis proposal–selection pipelines.
@@
Conventional RL explores only during training. Exceptions include works that explicitly optimize maximum entropy objectives (haarnoja2018soft) to learn stochastic policies, primarily to accelerate training convergence in settings where evaluation remains in-distribution, such as robotic control. Bayes-Adaptive RL (bellman1959adaptive; duff2002optimal; guez2012efficient; ghavamzadeh2015bayesian; lidayan2025bamdp; liu2023reason) has been studied to pursue the optimal exploration-exploitation trade-off in uncertain environments. When the true MDP identity is latent and must be inferred from interaction (states, actions, rewards), Bayesian RL connects naturally to Partially Observable MDPs (duff2001monte; ghosh2021generalization). Exact solutions of the Bayesian RL objective are often intractable, promoting the development of approximate methods (guez2012efficient; arumugam2022planning; chen2024bayes; arumugam2024satisficing). In our work, we adopt a policy gradient that operates over candidate answers, which differs from ghosh2022offline that leverages value ensembles in offline RL and qiu2025bayesian that applies SFT on an oracle Bayesian model’s outputs.
@@ 3 Problem Formulation
LLM Reasoning.
A finite-horizon MDP $\mathcal{M}$ can be defined by the state space $\mathcal{S}$, action space $\mathcal{A}$, horizon $T$, and reward function $r_{\mathcal{M}}(s,a)$, where $s\in\mathcal{S}$ and $a\in\mathcal{A}$. Here, the initial state $s_{0}$ is the prompt, and the action $a_{t}$ is the $t$-th step of the CoT, which can be either separated by special tokens (wang2023math) or defined as a fixed length of reasoning tokens (luo2024improve). We adopt the latter definition due to its simplicity. The state transition is deterministic by appending the new reasoning step, i.e., $s_{t+1}=s_{t}+a_{t}$.
We are interested in a training-deployment procedure. During training, the agent only receives information about the true MDP $\mathcal{M}^{*}\coloneqq(\mathcal{S},\mathcal{A},r,T)$ via the input training data $\mathcal{D}$. During deployment, we freeze the learned policy and deploy it in $\mathcal{M}^{*}$. Notably, $\mathcal{D}$ does not uniquely identify $\mathcal{M}^{*}$, inducing epistemic uncertainty about the MDP. Formally, $\mathcal{D}$ and a prior distribution over MDPs $p(\mathcal{M})$ define a posterior distribution over MDPs, $p(\mathcal{M}|\mathcal{D})\propto p(\mathcal{M})p(\mathcal{\mathcal{D}|\mathcal{M}})$. The objective for policy $\pi$ is thus to maximize return in expectation over MDPs from the posterior $p(\mathcal{M}|\mathcal{D})$:
$\mathcal{J}(\pi)=\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M}\mid\mathcal{D})}\Bigl[\mathcal{J}_{\mathcal{M}}(\pi)\Bigr],\,\,\,\,\text{where }\mathcal{J}_{\mathcal{M}}(\pi)=\mathbb{E}_{s_{0},\pi}\Biggl[\sum_{t=0}^{T-1}r_{\mathcal{M}}(s_{t},a_{t})\Biggr].$ (3.1)
This $\mathcal{J}(\pi)$ objective is designed to, for any data $\mathcal{D}$, train a policy $\pi$ that maximizes its Bayes-expected return in the unknown true environment during deployment, i.e., its performance averaged over MDPs drawn from the posterior $p(\mathcal{M}|\mathcal{D})$.
Progress Reward.
Prior work (uesato2022solving; guo2025deepseek) employs an outcome-level reward $\text{verifier}(s_{T},y_{s_{0}}^{*})$ for the true MDP $\mathcal{M}^{*}$, which uses a verifier to perform a regular expression match (either 0 or $1$) between $s_{T}$ and the ground-truth answer $y_{s_{0}}^{*}$ corresponding to the prompt $s_{0}$. We extend this sparse-reward setting by incorporating a progress reward (setlur2024rewarding; qu2025optimizing), which quantifies the increase of the model’s probability of outputting $y_{s_{0}}^{*}$ after appending $a$ at the CoT $s$, i.e., for $0\leq t\leq T-1$:
$r(s_{t},a_{t})=\pi_{\phi}(y_{s_{0}}^{*}\mid s_{t}+a_{t}+\text{</think>})-\pi_{\phi}(y_{s_{0}}^{*}\mid s_{t}+\text{</think>}),$ (3.2)
@@
$Q^{\pi}_{\mathcal{M}}(h_{t},a_{t})$ $=\mathbb{E}_{\pi}\Biggl[\sum_{t^{\prime}=t}^{T-1}r_{\mathcal{M}}(s_{t^{\prime}},a_{t^{\prime}})+\text{verifier}(s_{T},y_{s_{0}}^{*})\Biggr],$ (3.3)
where $h_{t}=(s_{0},a_{0},r_{0},\ldots,s_{t})$ is the history and we define $s(h_{t})=s_{t}$ as the final state $s_{t}$ from $h_{t}$.
Figure 1: An example of reflective reasoning trajectories.
@@
Informally, we define reflective exploration as the pattern in which the LLM revisits a prior state after an exploratory step to take different actions at that state. Specifically, a language reflection step such as “Let’s reconsider the geometric relationship” corresponds to revisiting and retrying. We illustrate this using a binary search tree as in the right figure: for the trajectory $s_{0}\shortrightarrow s_{1}\shortrightarrow s_{2}\shortrightarrow s_{1}\shortrightarrow s_{3}$, $s_{1}\shortrightarrow s_{2}$ is an exploration step, $s_{2}\shortrightarrow s_{1}$ is a reflective step that signals the strategy switch from $s_{2}$ to $s_{3}$, and the “geometric relationship” in the above example originates from $s_{1}$. We formalize this behavior through the lens of adaptive policies.
Definition 3.1 (Uncertainty-Adaptive Policy).
Define the belief as the posterior over MDPs $b(h_{t})(\mathcal{M})\coloneqq p(\mathcal{M}\mid h_{t},\mathcal{D})$. A policy $\pi$ is uncertainty-adaptive if there exists a measurable mapping $\mu:\mathcal{S}\times\mathcal{P}(\mathcal{M})\to\Delta(\mathcal{A})$ such that for all $h_{t}$ and $a_{t}$, $\pi(a_{t}\mid h_{t})=\mu(a_{t}\mid s(h_{t}),b(h_{t}))$. If in addition, there exists $h_{t}$ and $h_{t}^{\prime}$ such that $s(h_{t})=s(h_{t}^{\prime})$ but $\mu(a_{t}\mid s(h_{t}),b(h_{t}))\neq\mu(a_{t}\mid s(h_{t}^{\prime}),b(h_{t}^{\prime}))$, then $\pi$ is strictly uncertainty-adaptive, or equivalently, not Markovian.
From this perspective, reflective exploration is the behavioral signature of an uncertainty-adaptive policy: as the agent’s belief evolves, it rationally revisits previous states and generates a different action distribution. In contrast, a purely state-only policy that ignores this evolving belief would respond identically upon revisiting the same state. Formally,
Definition 3.2 (Reflective Exploration).
We say that a policy $\pi$ exhibits reflective exploration if there exist $t_{1}>t_{2}$ and $h_{t_{1}}$, $h_{t_{2}}$ such that
$\varphi(s(h_{t_{1}}))=\varphi(s(h_{t_{2}})),\qquad\pi(\cdot\mid h_{t_{1}})\neq\pi(\cdot\mid h_{t_{2}}),$
where $\varphi:\mathcal{S}\to\mathcal{Z}$ is a measurable mapping for which there exists $\vartheta:\mathcal{Z}\times\mathcal{H}_{\mathcal{R}}\to\Delta(\mathcal{A})$ satisfying for every $h_{t}$ that (1) $\vartheta(\cdot\mid\varphi(s_{t}),r_{0:t-1})=\pi(\cdot\mid h_{t})$; and (2) if there is some process $X_{t}$ and some $\upsilon$ such that $\pi(\cdot\mid h_{t})=\upsilon(\cdot\mid X_{t})$ then $X_{t}=\psi(\varphi(s_{t}))$ for some measurable $\psi$.
Intuitively, it describes an agent that returns to the same underlying latent state but, in light of what it has learned from past rewards, deliberately chooses a different strategy than it did before. The latent state representation $\varphi$ abstracts away superficial differences in the text state. The existence of $\vartheta$ and $\upsilon$, $\psi$ formalizes the latent factorization and state-minimality properties of $\varphi$, respectively.
4 The Necessity of Bayesian RL for Reflective Exploration
Conventional RL.
The RL objective for policy $\pi$ is $\mathcal{J}_{\mathcal{M}^{*}}(\pi)$ in (3.1). It is widely known that the optimal policy and value satisfy the Markov property by depending on $h_{t}$ only through the state $s_{t}$.
Theorem 4.1 (RL Admits Markovian Optimal Policy).
For any MDP, there exists a Markovian policy $\pi^{*}$, i.e., a policy that is not strictly uncertainty-adaptive, that maximizes the RL objective. Every policy that exhibits reflective exploration is not Markovian.
The result indicates that the RL objective admits Markovian policies to be optimal. Optimizing a Markovian policy, however, does not induce reflective exploration behaviors. Intuitively, action sequences whose effect is only to enrich the history $h_{t}$ (e.g., incorrect attempts followed by revisiting the same state) do not change the sufficient state representation $s_{t}$ and therefore cannot improve the value of an optimal Markovian policy.
Instead, in RL, exploration is only useful during training, in a trial-and-error manner with repeated episodes, to discover a return-maximizing action sequence. The learned policy is fully exploited at deployment time with no incentive for further exploration since no policy updates occur, which is why $\epsilon$-greedy often sets $\epsilon\approx 0$ when deployed (mnih2015human; hessel2018rainbow).
Bayesian RL.